OpenStudy (anonymous):

What is the solution set of y = x2 + 2x + 7 and y = x + 7? {(0, 7), (-1, 6)} {(0, 7), (-7, 0)} {(0, 7), (1, 8)} {(-2, 0), (4, 0)}

OpenStudy (rhystic):

there are a couple ways to do this. the easiast but longest way is just sub in teh coordinates and see if they satisfy the equation. for instance 0,7 satisfies both equations. -1,6 does not.

OpenStudy (campbell_st):

the easy way to solve this to to take the equations and set them equal to each other. This can happen because they both equal y. you are basically looking for the x values where the 2 curves intersect. so you have \[x^2 + 2x + 7 = x + 7\] start by collecting like terms... if done correctly you will have all the terms on the left and the right will be zero. then you solve by factoring to find the x values

OpenStudy (hockeychick23):

{(0, 7), (-1, 6)}

OpenStudy (anonymous):

x2+2x+7=x+7 x2+x+0=0 find delta=1*1-4(1*0) =1 find square root of delta=1 now x1=(-1+square root of delta=1)/2 x1=0 x2=(-1-1)/2 x2=-1 u will get two x cuz its a quadratic equation